### v1

# Are Time-Indexed Foundation Models the Future of Time Series Imputation?

Abstract
Foundation models for time series imputation remain largely unexplored. Recently, two such models, TabPFN-TS and MoTM, have emerged. These models share a common philosophy that places them within the family of time-indexed foundation models. This paper presents the first large-scale empirical study of these models for zero-shot imputation, which enables missing value recovery without retraining across a wide range of scenarios. We conduct extensive univariate experiments across 33 out-of-domain datasets ($\approx$ 1.3M imputation windows) and evaluate their ability to integrate covariates at inference time to improve accuracy without fine-tuning. Our results demonstrate that time-indexed foundation models are a powerful and practical step toward achieving general-purpose, zero-shot imputation for real-world time series.
00footnotetext: * Equal contribution
1 Introduction
Real-world time series from domains such as healthcare, industry, and climate science are often irregularly sampled or incomplete due to sensor failures and decentralized data collection (Schulz & Stattegger, 1997; Clark & Bjørnstad, 2004). Reliable imputation is thus a critical first step toward downstream tasks like forecasting, classification, or anomaly detection. Yet, while recent deep learning methods have advanced imputation performance (Cao et al., 2018; Du et al., 2023a), they typically lack robustness to distribution shifts and fail to generalize to out-of-domain data.
Recently, zero-shot forecasting models have emerged in the time series community, enabling inference on unseen datasets without retraining. This shift has given rise to time series foundation models, offering key benefits: (i) a single deployable model across diverse domains, (ii) strong performance on new datasets, often exceeding supervised baselines and (iii) emerging capabilities beyond simple memorization. While forecasting-oriented foundation models are now relatively well-studied (Auer et al., 2025b; Das et al., 2024; Woo et al., 2024; Ansari et al., 2024), imputation-focused counterparts remain scarce. Zero-shot imputation in out-of-domain settings is particularly challenging due to heterogeneous sampling rates, diverse missingness patterns, unaligned or irregular time series, and the potential presence of covariates whose predictive value is often underexploited.
A promising direction to overcome these challenges lies in continuous-time modeling, which learns a contextual representation $H(t)$ at each timestamp $t$, thereby casting imputation as a regression problem. Within this paradigm, time-indexed foundation models such as TabPFN-TS (Hoo et al., 2025) and MoTM (Le Naour et al., 2025) have recently been introduced, enabling zero-shot imputation through continuous-time representations. These models offer strong out-of-domain generalization without requiring retraining, making them particularly attractive for real-world applications where training data and computational resources are limited.
In this paper, we conduct the first extensive study of time-indexed foundation models for time series imputation, and summarize our key contributions as follows:
  • Extensive univariate evaluation. We evaluate TabPFN-TS and MoTM across 33 out-of-domain univariate datasets (covering roughly 1.3M windows to impute), benchmarking them against a wide range of baselines. TabPFN-TS yields the highest overall performance with a notable margin, whereas MoTM also surpasses all supervised and local baselines but remains behind TabPFN-TS.
  • Evaluating covariate integration without retraining. On three complex datasets, we show that both foundation models can seamlessly incorporate additional covariates at inference time, drastically improving imputation accuracy without any covariate-specific pretraining.
  • Limitations and discussion. We analyze the practical constraints of these approaches and identify scenarios where they are most effective. We also discuss potential directions toward more efficient and generalizable foundation models for time series imputation.
2 Considered Imputation Baselines for the Benchmark
This section presents the imputation baselines considered in the benchmark, covering a spectrum of approaches from simple statistical heuristics to modern foundation models. In Section˜2.1 we present local and supervised models, which, respectively, require dataset-specific training or rely on handcrafted rules; and in Section˜2.2 we present time-indexed foundation models, which generalize across datasets in a zero-shot manner without retraining. Further information on the models and their implementations can be found in Appendix˜A.
2.1 Local Imputers and Supervised Models
Local Imputers.
Within the benchmark, local imputation methods serve as fundamental baselines due to their simplicity, interpretability, and low computational cost. These approaches estimate missing values using only neighboring observations or straightforward statistical rules. Representative techniques include Linear Interpolation, which connects adjacent observations under a constant-rate assumption; Last Observation Carried Forward (LOCF), which propagates the most recent available value; and the Seasonal Naive method, which repeats the last observed value of a given periodicity (e.g., daily or weekly). While these methods perform adequately for small and isolated gaps, they generally fail to capture long-term dependencies, seasonal structures, or nonlinear dynamics commonly observed in real-world time series.
Supervised Models for Imputation.
Supervised models constitute the conventional approach for tackling complex imputation tasks, where models are trained end-to-end on specific datasets and evaluated on held-out test sets. These task-specific models — such as SAITS (Du et al., 2023a), BRITS (Cao et al., 2018), CSDI (Tashiro et al., 2021), TimesNet (Wu et al., 2023), or TimeMixer++ (Wang et al., 2025) — are typically deep learning architectures based on recurrent networks, attention mechanisms, or diffusion processes. Their main advantage lies in their ability to capture intricate temporal dependencies and model complex data distributions through direct optimization on the target dataset. However, their reliance on large training datasets often limits their generalization ability in zero-shot or cross-dataset scenarios, requiring retraining on each new task.
2.2 Time-Indexed Foundation Models for Imputation
The emergence of foundation models marks a paradigm shift toward general-purpose approaches for time series analysis. These models are characterized by their zero-shot capability: rather than being fine-tuned for the imputation task, they directly apply their pre-acquired knowledge to new data.
Models such as MoTM (Le Naour et al., 2025) and TabPFN-TS (Hoo et al., 2025) differ from conventional foundation models for time series forecasting, which often rely on patch-based attention architectures or extended LSTM variants (e.g., xLSTM in TiReX (Auer et al., 2025b)). Patch-based forecasters are trained to predict ground truth values over a given horizon conditioned on sequences of dense fully-observed contexts. However, such models do not handle the diverse missingness patterns of irregular time series inherent to the imputation task. On the other hand, both MoTM and TabPFN-TS adopt a continuous-time modeling design that naturally generalizes to unobserved timestamps and allows, at inference time, to: (i) handle irregular or unaligned time series, (ii) operate across different sampling rates, (iii) impute arbitrarily missing regions, and (iv) integrate covariates through concatenation with contextual representations.
In essence, these two time-indexed models learn a contextual representation $H(t)$ at every timestamp $t$. A regressor $r_{\theta}(\cdot)$ is then applied to map $H(t)$ to the observed time series value $x(t)$. Yet, despite their conceptual proximity, both models differ substantially in their architectural design. Below we describe both methods.
MoTM (Mixture of TimeFlow Models).
MoTM (Le Naour et al., 2025) extends the continuous-time modeling paradigm by leveraging a sophisticated feature extraction mechanism inherited from the TimeFlow architecture (Le Naour et al., 2024). Its core principle is to represent any time series through a pre-trained basis of modulated Implicit Neural Representations (INRs).
(i) Representation Learning via Modulated INR Basis. Specifically, MoTM does not learn a single function for the time series but rather a basis of $K$ distinct INRs. Each INR is a small neural network, parameterized by a hypernetwork (Dupont et al., 2022), that maps a continuous time coordinate $t$ to a feature vector. These basis functions are "modulated" in the sense that their parameters are dynamically generated for each new window, allowing them to capture a wide range of temporal patterns (e.g., trends, seasonalities, high-frequency oscillations) without being restricted to predefined frequencies like Fourier features. For any given timestamp $t$, the rich contextual representation $H(t)$ is formed by concatenating the outputs of all $K$ basis INRs evaluated at that time.
(ii) In-Context Imputation via Local Regression. The key mechanism of MoTM for imputation lies in its local, in-context fitting procedure. Given a time series with missing values, MoTM first considers a context window of observed points. It then fits a simple ridge regressor to learn the linear mapping from the high-dimensional representations $H(t)$ of these observed points to their actual values $x(t)$. This local regressor, fitted specifically on the available context, is finally used to predict the values at any missing timestamp $t_{\text{miss}}$ by applying it to the corresponding representation $H(t_{\text{miss}})$.
This framework naturally extends to: $\bullet$ Covariates integration with no retraining. Assuming full observation, additional contextual information available at timestamp $t$ are simply stacked to the target contextual representation $H(t)$. Ridge adaptation proceeds then in the same way as in the univariate setting, leaving the pretrained basis of INRs unchanged. $\bullet$ Quantification of uncertainty. By replacing the ridge regressor with a quantile regressor, MoTM can generate quantile predictions to produce uncertainty quantification intervals around the imputed values.
TabPFN-TS.
Hoo et al. (2025) apply the continuous-time modeling philosophy by reframing the time series imputation task as a standard tabular regression problem. This allows the direct application of the powerful, pre-trained TabPFN model (Hollmann et al., 2025) for zero-shot time series analysis. The model’s design philosophy is conceptually inverse to that of MoTM: it pairs a simple, handcrafted feature representation with a highly expressive regression model.
(i) Handcrafted Temporal Representations. Unlike MoTM’s learned representations, TabPFN-TS employs a straightforward feature engineering approach. The contextual representation $H(t)$ for each timestamp $t$ is constructed by combining the normalized time index itself with a set of pre-defined Fourier basis functions (i.e., sine and cosine pairs). These Fourier features are chosen to capture key seasonalities expected in the data (e.g., daily, weekly). This method results in a simple, fixed feature set that explicitly encodes temporal position and periodicity, serving as the input for the regression model (see Section˜A.3.2).
(ii) Imputation via In-Context Learning with TabPFN. The core expressive power of TabPFN-TS resides in its regressor, the TabPFN model. TabPFN is a large transformer-based architecture pre-trained on hundreds of millions of synthetically generated tabular regression tasks. Its defining characteristic is in-context learning: at inference time, it ingests a set of observed data points—pairs of temporal features and their corresponding values, $(H(t_{obs}),x(t_{obs}))$ as a single "prompt." The model processes this entire context within its attention layers to infer the underlying functional relationship between the features and the series values. It then applies this inferred function to predict the values $x(t_{\text{miss}})$ for the query features $H(t_{\text{miss}})$ of missing timestamps, all within a single forward pass and without any gradient-based fine-tuning.
This framework also naturally supports the Integration of covariates with no retraining and Uncertainty quantification. In particular, TabPFN inherently models uncertainty by returning distribution over outputs.
3 Experiments
We design our experimental study to assess two key aspects: (i) zero-shot generalization across out-of-domain datasets and (ii) the ability to incorporate auxiliary covariates without retraining. Thus, experiments are organized into two main parts: a large-scale univariate benchmark covering 33 datasets (Section˜3.1), and a focused covariate integration study on three datasets (Section˜3.2).
3.1 Univariate Benchmark: Out-of-Domain Zero-Shot Imputation
In this section, we evaluate the out-of-domain (OoD) zero-shot performance of TabPFN-TS and MoTM across 33 diverse real-world datasets. These datasets cover a wide range of sampling rates (5min, 10min, 15min, 30min, 1h) and exhibit heterogeneous temporal patterns and seasonalities. They originate from open-source collections spanning multiple domains, including climate, energy, traffic, etc. Most are drawn from LOTSA (Woo et al., 2024) and GIFT-eval (Aksu et al., 2024), with strict safeguards to prevent leakage from the three pretraining datasets of MoTM. For all details on the datasets please refer to Section˜B.1. In total, the evaluation involves more than 1.3M incomplete windows.
Protocol and Baselines.
All datasets are split chronologically intro train, validation and test fractions with respective ratios specified in Section˜B.1. The test split is divided into four-week segments. For each window, we generate four distinct missing data scenarios, by randomly removing: either (i) 50% and (ii) 70% of the observations (Pointwise scenarios); or (iii) two and (iv) four entire days (Block scenarios). Note that only the supervised baselines (denoted as Task Specific Models) benefit from the train and validation sets. The benchmark includes all the methods described in Section˜2. Implementation details, hyperparameters, and method-specific configurations are provided in Appendix˜A.
3.1.1 Main Results
The aggregated results of our univariate benchmark are shown in Figure˜1, summarizing the mean normalized MAE across all 33 out-of-domain datasets. Each bar represents the overall imputation accuracy of a model averaged over the four missingness regimes. Models are grouped into three categories reflecting their underlying paradigm: (i) localmethods, which impute from within each series without any learned representation; (ii) task-specificmodels, trained in a supervised manner on each dataset; and (iii) foundation modelsevaluated in a fully zero-shot setting. To complement these aggregated results, we provide in Figure˜2 a critical difference (CD) diagram (Ismail Fawaz et al., 2019) summarizing the average ranks of all models across datasets and missingness scenarios.
Figure 1: Univariate Benchmark on Out-of-Domain datasets, reported results are z-normalized MAEs.
Results.
As shown in Figures˜1 and 2, three consistent trends emerge from the benchmark.
(i) Time-index foundation models lead the benchmark. Both TabPFN-TS and MoTM achieve the lowest aggregated normalized errors and the best average ranks, substantially outperforming all supervised and local baselines despite being evaluated in a fully zero-shot setting. The CD diagram further confirms that these two models are statistically superior compared to all others, except MoTM compared to SAITS. Among them, TabPFN-TS attains the top performance (MAE = 0.293, avg. rank = 1.35), suggesting that large-scale synthetic pretrained regressor combined with explicit temporal encodings confers a measurable advantage over the Ridge on top of learned continuous representations.
(ii) Task-specific models show limited robustness. While SAITS achieves competitive accuracy (0.386, avg. rank = 3.56), other supervised approaches such as BRITS, CSDI, or TimesNet lag behind, sometimes performing worse than simple local heuristics (e.g., Seasonal Naive, LOCF). This mixed behavior highlights the limited generalization capacity of fully supervised models, which tend to overfit dataset-specific temporal dynamics — particularly when training data is scarce.
(iii) Local baselines remain resilient. Classical approaches leveraging temporal priors still deliver reasonable performance in heterogeneous settings, as reflected in both their MAE and ranking consistency. However, the pronounced gap separating them from foundation models clearly illustrates the benefits of pretraining time-indexed models, which provide both accuracy and adaptability without retraining.
Figure 2: Critical difference diagram over all 33 univariate OOD zero-shot imputation tasks.
Overall, the aggregated metrics and the rank-based analysis provide clear evidence that foundation models — particularly TabPFN-TS — deliver the most consistent and robust performance for zero-shot time-series imputation across diverse domains. MoTM also achieves an honorable level of accuracy despite relying on a comparatively simple regressor, underscoring the effectiveness of its pretrained time-indexed representations. Complete benchmark results are reported in Section˜C.1, while additional experiments on datasets with lower sampling rates (daily and weekly) are provided in Section˜C.4.
3.1.2 Uncertainty Quantification Results
Beyond pointwise accuracy, it is important to assess how well models capture predictive uncertainty. Both TabPFN-TS and MoTM natively support quantile estimation, allowing them to produce calibrated uncertainty bounds around each imputed value. Among the baselines, only CSDI provides comparable quantile predictions.
We report in Table˜1 the average Weighted Quantile Loss (WQL) (see Section˜C.3 for loss definition) across eleven representative datasets (full results and more details in Section˜C.3.2). This metric evaluates both imputation accuracy and quantile calibration. The WQL is computed over nine quantile levels, from 0.1 to 0.9, providing a comprehensive measure of the models’ probabilistic consistency across the predictions.
Table 1: Weighted Quantile Loss (WQL) average scores on eleven representative univariate datasets.
TabPFN-TS MoTM CSDI SAITS-Q
WQL 0.241 0.316 0.453 0.476
Results.
Quantitatively, TabPFN-TS achieves the lowest overall WQL score, followed by MoTM and CSDI. SAITS-Q, a variant of SAITS where one model is trained independently per quantile level, obtains the highest WQL. This suggests that foundation-based models such as TabPFN-TS and MoTM not only reconstruct missing values more accurately but also provide better-calibrated uncertainty estimates.
Qualitative examples in Figure˜3 confirm these trends: both foundation models adjust their uncertainty bands (5–95 and 25–75 quantile ranges) to local signal variability. TabPFN-TS produces sharper, high-fidelity reconstructions that closely follow the ground truth, while MoTM yields smooth yet robust imputations with stable and well-calibrated uncertainty envelopes. Additional examples are provided in Section˜C.3.3.
(a) TabPFN on JenaWeatherH dataset
(b) MoTM on Hog dataset
Figure 3: Qualitative quantile results in the 70% missing values scenario (Pointwise 2).
3.2 Integration of Covariates in Zero-Shot OoD Imputation
For a model to be practically deployable in real-world applications, it must be capable of incorporating covariates, which often enhance predictive accuracy by providing additional contextual information. This requirement is recognized as one of the major challenges for foundation models in time series forecasting as well (Auer et al., 2025a). In this section, we investigate how TabPFN-TS and MoTM handle covariates in a zero-shot setting. We conduct experiments on three datasets where the dependence of the time series values on covariates ranges from weak to critical, in order to evaluate how efficiently they integrate this additional information into their imputations.
Datasets, Baselines, and Protocol.
We evaluate the models on imputation tasks across three datasets: $\bullet$ PV-France; aggregated photovoltaic (PV) production curves with the associated global solar irradiance as a covariate, $\bullet$ Wind-France; aggregated wind production curves with wind speed as a covariate, and $\bullet$ Load-France; aggregated electricity consumption curves with temperature as a covariate. Further details on the datasets and preprocessing are provided in Section˜B.2.
The train / validation / test splits follow the same protocol as in Section˜3.1. A key difference is that at inference, the covariate is always assumed to be fully observed. We compare the following methods: TabPFN-TS and MoTM both with or without covariates, a ridge regression using only the covariate to predict the target variable, and SAITS. Results are showcased in Table˜2.
Table 2: Complete MAE results on datasets with covariates. The best score for each setting is shown in bold, second-best is underlined.
TabPFN-TS MoTM Other baselines
Dataset Setting
TabPFN-TS
(W/ Covar)
TabPFN-TS
(W/o Covar)
MoTM
(W/ Covar)
MoTM
(W/o Covar)
Ridge
on Covar
SAITS
Multivar
PV-France Pointwise 1 0.045 0.109 0.054 0.102 0.115 0.390
Pointwise 2 0.051 0.160 0.069 0.131 0.123 0.443
Blocks 1 0.052 0.175 0.086 0.191 0.104 0.496
Blocks 2 0.049 0.190 0.086 0.179 0.106 0.485
Wind-France Pointwise 1 0.101 0.098 0.128 0.186 0.318 0.359
Pointwise 2 0.138 0.153 0.161 0.244 0.322 0.408
Blocks 1 0.275 0.470 0.335 0.600 0.321 0.590
Blocks 2 0.248 0.470 0.317 0.594 0.335 0.581
Load-France Pointwise 1 0.037 0.037 0.138 0.146 0.667 0.292
Pointwise 2 0.056 0.059 0.158 0.164 0.667 0.321
Blocks 1 0.143 0.146 0.236 0.243 0.669 0.490
Blocks 2 0.170 0.178 0.262 0.270 0.674 0.498
Integrating covariate improvement / 31.54% / 31.03% / /
Results.
As shown in Table˜2, incorporating covariates generally improves performance, both for TabPFN-TS and MoTM. Substantial gains are observed on datasets for which the covariate strongly informs about the target variable (global irradiance for PV power production, wind speed for wind power production). Under these scenarios and particularly in the challenging block settings, incorporating covariates allows MoTM to outperform the impressive univariate performances of TabPFN-TS. This emphasizes how fundamental it is to enhance pretrained univariate foundation models with additional contextual information for them to be deployable in real-world applications. On the other hand, weaker relationships, such as between national-level electricity demand and average temperature, yield marginal to no improvement. Overall, TabPFN-TS (With Covar) achieves the best results in most scenarios, confirming its strong capacity to integrate heterogeneous contextual inputs for robust zero-shot imputation.
3.2.1 Qualitative Covariates Results
To better visualize the impact of incorporating covariates, we present qualitative results for both TabPFN-TS and MoTM. Figure˜4 and Figure˜5 show examples from the Wind-France dataset, illustrating how the inclusion of covariate information helps each model reconstruct the four one-day missing blocks more accurately.
Figure 4: Wind-France dataset. TabPFN-TS qualitative results with and without covariates in the four one-day missing blocks scenario.
Figure 5: Wind-France dataset. MoTM qualitative results with and without covariates in the four one-day missing blocks scenario.
Results.
As illustrated in Figures 4 and 5, incorporating covariates visually improves the reconstructions for both TabPFN-TS and MoTM. The covariate-enhanced versions capture missing intervals more smoothly and better follow short-term temporal variations within the four one-day gaps. We observe that the interpolations of TabPFN-TS align more closely with the ground truth and preserve sharper transitions across missing regions. In contrast, if MoTM consistently benefits from covariate information, its reconstructions sometimes struggle to capture sudden changes with great accuracy. More covariate plots are showcased in Appendix˜D, confirming that the most visible gains are observed on the Wind-France and PV-France datasets, while the impact remains limited for Load-France.
4 Practical Considerations
While the previous sections establish the strong zero-shot performance of foundation models, it is equally important to assess their practical limitations and behavioral consistency under different conditions. This section provides complementary analyses covering two key aspects: (i) the computational efficiency of the proposed models across datasets of varying scales, and (ii) their robustness to different missingness patterns.
4.1 Computational Cost
In this section, we compare the computational efficiency of the proposed models across datasets of varying sizes in Figure˜6. The reported times correspond solely to inference for TabPFN-TS, MoTM, and Linear, as these methods operate in a zero-shot manner. For the supervised baseline SAITS, we report the total time including both training and inference. The analysis spans four representative datasets — ranging from small (Covid19 Energy) to large-scale (Traffic) — to illustrate the scalability and practical trade-offs between methods. For each dataset, the reported MAE results are averaged over the four missing-data scenarios. All experiments are carried out on a single NVIDIA H100-80G GPU.
Figure 6: Compute time (log-scale) versus MAE across datasets of increasing size.
Results.
As shown in Figure˜6, TabPFN-TS consistently achieves the lowest MAE but at the cost of substantially higher inference time, especially on large datasets. In contrast, MoTM offers a favorable balance between accuracy and efficiency, being up to two orders of magnitude faster while maintaining competitive performance. SAITS attains moderate error levels but requires significant computational overhead due to its supervised training phase. Finally, Linear is the fastest overall but shows markedly poorer accuracy, particularly on more complex datasets. These results highlight the trade-off between zero-shot generalization and computational cost, emphasizing MoTM as a scalable alternative to TabPFN-TS.
4.2 Breakdown by Missing Patterns
A detailed breakdown of results across the four missingness regimes is provided in Section˜C.2. The analysis highlights a clear dependence of model performance on the structure of the missing values. For sparse pointwise removal (50–70% pointwise removal), Linear Interpolation remains highly competitive, occasionally matching foundation models—indicating that simple local continuity assumptions suffice when gaps are isolated. However, performance rapidly degrades under structured, block-wise missingness (two and four one-day gaps), where both TabPFN-TS and MoTM maintain stable accuracy and clearly dominate all baselines. Overall, these results confirm the robustness of foundation models to temporally correlated gaps, while emphasizing that local methods remain effective only in low-missingness, unstructured regimes.
5 Conclusion and Discussion
Our experiments demonstrate that TabPFN-TS achieves very strong zero-shot performance, both in univariate imputation and when integrating covariates. Nevertheless, its inference times remain a significant limitation for some real-world deployment scenarios. MoTM also delivers strong zero-shot performance, outperforming all supervised baselines on the univariate benchmark. Similarly to TabPFN-TS, its ability at leveraging additional contextual information is remarkable. However, it is generally less accurate than TabPFN-TS, although it offers substantially faster inference.
These two highly flexible foundation models represent a significant step forward toward “off-the-shelf” zero-shot imputation solutions applicable across a wide range of domains. A promising avenue to combine performance and efficiency would be to build a more powerful regressor on top of the modulated INR features used by MoTM, replacing the current ridge regressor with a model trained via in-context learning, potentially merging the strengths of both approaches.
Appendix A Baselines details and implementation
In this section, we present the imputers included in our experimental evaluation and detail their implementation. As introduced in Section 2, our selection spans a broad spectrum: from simple naïve baselines, through state-of-the-art supervised deep-learning models, to recent time-indexed foundation models. This diversity ensures coverage of both lightweight, assumption-free methods and advanced architectures capable of capturing long-range and cross-variable dependencies.
A.1 Local imputers
We considered three naïve baselines: a simple linear interpolation, a last observation carried forward, and a seasonal repeat. Note that each imputer operates independently on each incomplete segment and does not use any cross-variable information.
Linear.
This baseline imputes a missing value at timestamp $t$ by linearly interpolating between the closest observed neighbors surrounding the gap. Concretely, it uses the last observation before $t$ and the first observation after $t$ as anchors and fills the interior points on the straight line between them. If a leading gap has no past anchor, we adopt a next-observation-carried-backward (NOCB) fallback; if a trailing gap has no future anchor, we fall back to last-observation-carried-forward (LOCF).
LOCF.
This baseline imputes a missing value at timestamp $t$ by copying the most recent available past value (the last observation before $t$). If the series begins with a gap and no past value exists, we perform a single NOCB initialization by copying the first available future observation backward.
Seasonal Naive.
This baseline imputes a missing value at timestamp $t$ by using the observation from the previous seasonal period, i.e., the value at tSt-S. The seasonal period $S$ is pre-defined for each dataset based on its dominant frequency (e.g., daily or weekly). If the value at tSt-S is also missing, the method sequentially searches for an available observation at other seasonal timestamps (e.g., $t+S$, then t2St-2S, etc.). If this search fails to yield a value, the method falls back to a simple LOCF imputation.
A.2 Supervised imputers
We considered six supervised deep-learning baselines spanning complementary neural paradigms: recurrent, Transformers, and diffusion-based approaches. These include methods built specifically for imputation as well as recent multi-task time-series backbones (convolutional and token-mixing/Transformer hybrids) that tackle imputation via masked-reconstruction training. This diversity probes different inductive biases and offers a balanced accuracy/efficiency trade-off.
All models implementations are taken from the PyPOTS (Du et al., 2023b) Python toolbox and trained on fixed-length windows where a subset of observed entries is randomly masked; each model reconstructs these masks from the remaining context (values ++ binary observation mask). Inputs are z-score normalized per variable, training minimizes the mean absolute error (MAE) on masked positions only, model selection uses validation MSE with early stopping, and test metrics are computed on the 4 missing points scenario detailed in Section 3.1. Hyperparameter choices were inspired by the default settings recommended by each method’s authors in their original papers and/or public implementations. Note that we use the Adam optimizer (Kingma & Ba, 2015) for all models with a batch size of 64, train for at most 50 epochs, and apply early stopping with a patience of 5 epochs.
SAITS.
A Transformer imputer with two diagonally-masked self-attention (DMSA) blocks that jointly capture temporal and cross-feature dependencies; a learned gating mechanism combines both blocks to predict missing values efficiently (Du et al., 2023a).
Hyperparameters. We use $n_{\text{layers}}=2$, $d_{\text{model}}=256$, $n_{\text{heads}}=4$ with $d_{k}=d_{v}=64$, feed-forward size $d_{\text{ffn}}=128$ and dropout =0.1=0.1. This setting follows the authors’ recommended defaults and balances capacity with training stability.
BRITS.
A bidirectional RNN imputer with learned temporal decay that processes each window forward and backward, jointly estimating hidden states and missing values while enforcing consistency between directions (Cao et al., 2018).
Hyperparameters. A single-stack bidirectional GRU with $\text{rnn\_hidden\_size}=64$ and dropout =0.3=0.3. This mirrors common BRITS configurations and provides a compact baseline for irregular gaps.
CSDI.
A conditional diffusion imputer that models $p(\mathbf{x}_{\text{miss}}\mid\mathbf{x}_{\text{obs}})$ via score-based denoising; training adds Gaussian noise to targets and conditions the denoiser (Transformer/U-Net with time embeddings) on values and masks, enabling stochastic imputations at inference (Tashiro et al., 2021).
Hyperparameters. Transformer/U-Net backbone with $n_{\text{layers}}=4$, $n_{\text{heads}}=4$, $n_{\text{channels}}=64$; time, feature, and diffusion embeddings of sizes $128$, 3232, and $128$, respectively. We use $n_{\text{diffusion\_steps}}=50$, and a quadratic noise schedule type. These values track the public defaults for efficient training while preserving sampling quality.
TimesNet.
A multi-periodic convolutional model that folds each 1D series into 2D tensors along discovered periods and applies multi-scale 2D CNN blocks to capture intra- and cross-period structure; we train it for masked-value reconstruction (Wu et al., 2023).
Hyperparameters. We set $n_{\text{layers}}=2$, $d_{\text{model}}=128$, $d_{\text{ffn}}=256$, $n_{\text{kernels}}=6$, top_k=3=3, and dropout =0.1=0.1. This lightweight configuration keeps computation modest while retaining the multi-periodic 2D convolutional capacity.
TimeMixer++.
A token-mixing hybrid (Transformer/MLP) designed for multi-periodic patterns with lightweight mixing blocks and skip connections; adapted here to imputation by reconstructing randomly masked entries from the observed context (Wang et al., 2025).
Hyperparameters. We use $n_{\text{layers}}=3$, $d_{\text{model}}=64$, $d_{\text{ffn}}=128$, top_k=8=8, $n_{\text{kernels}}=6$, $n_{\text{heads}}=4$, with channel_mixing and channel_independence, dropout =0.1=0.1, and downsampling ($n_{\mathrm{downsampling\_layers}}=3$ and $n_{\mathrm{downsampling\_window}}=2$). This mirrors the authors’ small/efficient setting adapted for masked reconstruction.
TSLANet.
A lightweight convolutional model replacing self-attention with an Adaptive Spectral Block (Fourier-domain features with adaptive thresholding) and an Interactive Convolution Block for local–global mixing; trained for single-pass reconstruction of masked entries (Eldele et al., 2024).
Hyperparameters. Three layers ($\text{n\_layers}=3$) with patch_size =16=16, embedding dimension $d_{\text{embedding}}=256$, mask_ratio =0.4=0.4, and dropout =0.1=0.1 ; matching the default hyperparameters of the official codebase.
A.3 Time-Index Foundation models
In this section, we provide further architectural and implementation details for the two time-index foundation models evaluated in our experiments: MoTM and TabPFN-TS.
A.3.1 MoTM.
MoTM (Mixture of TimeFlow Models) extends the continuous-time modeling paradigm by leveraging a sophisticated feature extraction mechanism inherited from the TimeFlow architecture (Le Naour et al., 2024). Its core principle is to represent any time series through a pre-trained basis of modulated Implicit Neural Representations (INRs).
Representation Learning via Modulated INR Basis.
Specifically, MoTM does not learn a single function for the time series but rather a basis of $K$ distinct INRs. Each INR is a small neural network, parameterized by a hypernetwork (Dupont et al., 2022), that maps a continuous time coordinate $t$ to a feature vector. These basis functions are "modulated" in the sense that their parameters are dynamically generated for each new segment, allowing them to capture a wide range of temporal patterns (e.g., trends, seasonalities, high-frequency oscillations) without being restricted to predefined frequencies like Fourier features. For any given timestamp $t$, the rich contextual representation $H(t)$ is formed by concatenating the outputs of all $K$ basis INRs evaluated at that time.
In-Context Imputation via Local Regression.
The key mechanism of MoTM for imputation lies in its local, in-context fitting procedure. Given a time series with missing values, MoTM considers a context window of observed points. It then fits a simple ridge regressor to learn the linear mapping from the high-dimensional representations $H(t)$ of these observed points to their actual values $x(t)$. This local regressor, fitted specifically on the available context, is then used to predict the values at any missing timestamp $t_{\text{miss}}$ by applying it to the corresponding representation $H(t_{\text{miss}})$.
This framework naturally extends to:
  1. 1.
    Integrate covariates with no retraining. Assuming full observation, additional contextual information available at timestamp $t$ are simply stacked to the target contextual representation $H(t)$. Ridge adaptation proceeds then in the same way as in the univariate setting.
  2. 2.
    Uncertainty quantification. By replacing the ridge regressor with a quantile regressor, MoTM can generate quantile predictions to construct confidence intervals around the imputed values.
In essence, MoTM decouples the problem: (i) a powerful, pre-trained feature extractor (the INR basis) captures complex temporal dynamics, (ii) while a simple, locally-fitted linear model performs the final task-specific regression. This design allows the model to adapt to the specific structure of each new time series without requiring any fine-tuning of its core components.
Implementation Details.
Our implementation strictly adheres to the architecture and methodology described in the original paper (Le Naour et al., 2025). We utilize the three publicly available pre-trained TimeFlow (modulated INRs), which were respectively trained on the Electricity, Solar10T, and Spanish Temperatures datasets (further details on these datasets are available in Section˜B.1). For full reproducibility, both the source code and the pre-trained INR weights were obtained from the official repository: https://github.com/EtienneLnr/MoTM.
A.3.2 TabPFN-TS.
TabPFN-TS Hoo et al. (2025) applies the continuous-time modeling philosophy by reframing the time series imputation task as a standard tabular regression problem. This allows the direct application of the powerful, pre-trained TabPFN model (Hollmann et al., 2025) for zero-shot time series analysis. The model’s design philosophy is conceptually inverse to that of MoTM: it pairs a simple, handcrafted feature representation with a highly expressive regression model.
Handcrafted Temporal Representations.
Unlike MoTM’s learned representations, TabPFN-TS employs a straightforward feature engineering approach. The contextual representation $H(t)$ for each timestamp $t$ is constructed by combining the normalized time index itself with a set of pre-defined Fourier basis functions (i.e., sine and cosine pairs). These Fourier features are chosen to capture key seasonalities expected in the data (e.g., daily, weekly). This method results in a simple, fixed feature set that explicitly encodes temporal position and periodicity, serving as the input for the regression model.
Imputation via In-Context Learning with TabPFN.
The core expressive power of TabPFN-TS resides in its regressor, the TabPFN model. TabPFN is a large transformer-based architecture pre-trained on hundreds of millions of synthetically generated tabular regression tasks. Its defining characteristic is in-context learning: at inference time, it ingests a set of observed data points—pairs of temporal features and their corresponding values, $(H(t_{obs}),x(t_{obs}))$ as a single "prompt." The model processes this entire context within its attention layers to infer the underlying functional relationship between the features and the series values. It then applies this inferred function to predict the values $x(t_{\text{miss}})$ for the query features $H(t_{\text{miss}})$ of missing timestamps, all within a single forward pass and without any gradient-based fine-tuning.
This framework naturally supports: (i) Integratation of covariates with no retraining.Covariates available at a timestamp $t$ are simply concatenated to the temporal representation $H(t)$, expanding the feature set that TabPFN uses for its in-context learning. (ii) Uncertainty quantification.TabPFNinherently supports probabilistic predictions. By configuring it to predict specific quantiles instead of the mean, we can generate confidence intervals for the imputed values, as demonstrated in Figure˜8.
In essence, TabPFN-TS also decouples the problem, but with an inverted philosophy compared to MoTM: (1) a simple, handcrafted feature extractor creates temporal features, (2) while a highly expressive, pre-trained transformer model performs the complex in-context regression. This design leverages the generalization capability of a large pre-trained model to solve the imputation task.
Implementation Details.
For our experiments, we utilize the official TabPFNv2 implementation from the tabpfn Python package, which includes the pre-trained TabPFNv2 model described in (Hollmann et al., 2025). We employ a consistent feature representation, $H(t)$, across all datasets, defined as:
$H(t)=\left(t,\sin\left(\frac{2\pi t}{P_{\text{day}}}\right),\cos\left(\frac{2\pi t}{P_{\text{day}}}\right),\sin\left(\frac{2\pi t}{P_{\text{week}}}\right),\cos\left(\frac{2\pi t}{P_{\text{week}}}\right)\right)$
In this formulation, $t$ is the normalized time index, $P_{\text{day}}$ represents the number of timesteps (normalized) in a 24-hour period (e.g., 24 for hourly data, 48 for 30-minute data, 96 for 15-minute data etc.), and $P_{\text{week}}$ is set to $7\times P_{\text{day}}$ to capture weekly seasonality. For full reproducibility, the source code and pre-trained regressor are publicly available at https://github.com/PriorLabs/TabPFN/.
A.4 Excluded baselines
MOMENT.
We considered including MOMENT (Goswami et al., 2024), a large Transformer-based foundation model pre-trained on patched time series via a masked modeling objective. Initially, our intent was to evaluate it as a zero-shot imputation baseline alongside TabPFN-TS and MoTM. However, its performance in our benchmark was found to be substantially lower than that of the other methods. A closer review of the original paper clarifies this result: the authors primarily advocate for fine-tuning the model on specific downstream tasks to achieve optimal performance. Moreover, their reported imputation experiments are limited to very short missing segments, a scenario that differs significantly from those addressed in our work. Given its unsuitability for zero-shot time series imputation, we excluded it from our final comparative analysis.
Other deep learning supervised imputers.
The field of deep learning for supervised time series imputation is a rapidly evolving area of research. Our selection of baseline models aims to provide a representative sample of this landscape, including both supervised methods that have become standard benchmarks and more recent, state-of-the-art architectures. To ensure fair and reproducible comparisons, all selected deep learning supervised models are part of the PyPOTS library (Du et al., 2023b). This framework was essential for conducting our extensive benchmark.
Appendix B Datasets details
B.1 Univariate datasets
The complete list of datasets used in our univariate experiments is shown in Table˜3. 36 datasets were used in total: 3 for the pretraining of MoTM and the remaining 33 for the zero-shot evaluation of both TabPFN-TS and MoTM. Each dataset is split chronologically intro train, validation and test. Unless otherwise stated, the respective fractions are 0.7, 0.1 and 0.2. The test segments were then generated by applying a four-week sliding window, where at every step a random stride is drawn uniformly between 0.5 and 2 days. This procedure ensures that the inference samples are not aligned on any specific calendar information. A short description of each dataset is provided below. Those datasets marked with an "*" were curated to remove flat segments at inference to avoid biasing the evaluation towards trivial scenarios. Flat segments are caused e.g. by heterogeneous sensor operating dates within datasets or by filling long missing blocks with zeros (Emami et al., 2023).
Table 3: All datasets used in our univariate experiments and their key properties.
Dataset Release Domain MoTM Freq Num. Series Num. Test
Platform Use Series Length Segments
Electricity Zenodo Energy Train 1H 370 35064 122623
Solar-10T Zenodo Energy Train 10min 137 52560 9179
Spanish Temperatures Kaggle Climate Train 1H 5 35000 1090
BDG2-Bear LOTSA Energy Test 1H 91 17544 7522
BDG2-Rat LOTSA Energy Test 1H 280 17544 24915
Borealis LOTSA Energy Test 1H 15 7447 77
Covid19 Energy LOTSA Energy Test 1H 1 31912 195
GFC12 Load LOTSA Energy Test 1H 20 39414 4960
Hog LOTSA Energy Test 1H 24 17544 2310
Ideal LOTSA Energy Test 1H 217 16167 156
PDB LOTSA Energy Test 1H 1 17520 96
KDD Cup2022 LOTSA Energy Test 10min 134 35279 2546
ERA5 geopotential LOTSA Climate Test 1H 500 8736 19000
ERA5 humidity LOTSA Climate Test 1H 500 8736 19000
ERA5 temperature LOTSA Climate Test 1H 500 8736 19000
ERA5 wind speed LOTSA Climate Test 1H 500 8736 19000
Oikolab Weather LOTSA Climate Test 1H 8 100057 5288
Pedestrian Counts LOTSA Transport Test 1H 66 96400 7733
Traffic LOTSA Transport Test 1H 861 17544 83479
PEMS BAY LOTSA Transport Test 5min 325 52128 2275
PEMS 03 LOTSA Transport Test 5min 358 26208 358
SHMETRO LOTSA Transport Test 15min 576 8809 576
ETT1-15T GIFT-eval Energy Test 15min 7 69680 1050
ETT1-1H GIFT-eval Energy Test 1H 7 17420 1092
ETT2-15T GIFT-eval Energy Test 15min 7 69680 1050
ETT2-1H GIFT-eval Energy Test 1H 7 17420 1092
Solar-1H GIFT-eval Energy Test 1H 137 8760 8768
Jena Weather 10T GIFT-eval Climate Test 10min 21 52704 1428
Jena Weather 1H GIFT-eval Climate Test 1H 21 8784 1344
Loop Seattle 5T GIFT-eval Transport Test 5min 323 105120 21964
Loop Seattle 1H GIFT-eval Transport Test 1H 323 8760 20672
MDense GIFT-eval Transport Test 1H 30 17520 4710
Enedis LDM Small Zenodo Energy Test 30min 500 17424 20500
London Smart Meters Small Chronos Energy Test 30min 500 22000 25779
Spanish Energy Kaggle Energy Test 1H 9 35064 1962
Weather Informer Climate Test 1H 11 35064 2398
Energy domain
BDG2-Bear*, BDG2-Rat* and Hog* are the energy demands of commercial buildings in the US in 2016 - 2017. Sourced by the BuildingsBench library (Emami et al., 2023) from the Building Data Genome 2 (BDG2) project (Miller et al., 2020).
Borealis* and Ideal* contain the total electricity consumption of, respectively, 15 homes in Waterloo, Ontario, in 2011-2012 and 217 homes in Edinburgh, UK, between 2016-2018. Both datasets were released as part of the BuildingsBench dataset, and include a marginal amount of data preprocessing (including interpolation of missing values and outlier removal) (Emami et al., 2023).
Covid19 Energy is the aggregated electricity demand of an entire metropolitan area, from 2017 to 2021 (Farrokhabadi et al., 2022).
GFC12 Load is sourced from the Global Energy Forecasting Competition 2012 and contains a total of 20 aggregated load series (Wang et al., 2023).
PDB is a Kaggle dataset containing electricity demand and outdoor temperature data in 2013-2014. We omitted the temperature and kept only the electricity demand (Wang et al., 2023).
Electricity contains the hourly-aggregated electricity consumption of 370 households in Portugal in 2011-2014 (Trindade, 2015).
KDD Cup 2022 is a dataset for the Spatial Dynamic Wind Power Forecasting Challenge hosted at KDD in 2022 (Zhou et al., 2022). It contains the wind power data of 134 wind turbines from a wind farm over half a year. We kept the wind power generation variable as the target variable and omitted extra covariates (wind speed and direction, temperature, etc.). The train / validation / test split is 0.65 / 0.15 / 0.2.
Solar contains the synthetic power production of 137 photovoltaic power plants in Alabama in 2006. Dataset sourced by Lai et al. (2018) using simulations from NREL’s Solar Power Data for Integration Studies.
ETT1 and ETT2 respectively contain measurements of oil temperature of two electrical transformers in China, as well as six additional covariates. We used these 7 variables in our experiments, handling them with channel independence. The datasets were collected and published by Zhou et al. (2021).
London Smart Meters Small is the half-hourly energy consumption of 5561 households in the UK between 2011 and 2014. Data sourced by Godahewa et al. (2020) from https://data.london.gov.uk/dataset/smartmeter-energy-use-data-in-london-households. We kept a random subset of 500 samples in our experiments.
Enedis LDM Small is a dataset of 10k one-year individual electricity consumptions generated by a latent diffusion model at a 30-min sampling rate and representative of thermo-sensitive French households (Nabil et al., 2025). We kept a random subset of 500 samples in our experiments.
Spanish Energy is a Kaggle dataset containing the electricity (i) consumption and (ii) production for Spain from 2015 to 2018. We used the total load demand as well as electricity generation of eight energy sources (biomass, fossil gas, fossil hard coal, solar, onshore wind and three technologies of hydropower). We kept these nine variables in our experiments, handling them with channel independence. Data were obtained from https://www.kaggle.com/datasets/nicholasjhana/energy-consumption-generation-prices-and-weather.
Climate domain
Spanish Temperatures contains the hourly temperature measurements for the five largest cities in Spain, from 2015 to 2018. Data were obtained from https://www.kaggle.com/datasets/nicholasjhana/energy-consumption-generation-prices-and-weather.
ERA5 is part of the ClimateLearn library, which provides historical worldwide time series of various climate (atmosphere and land-surface) variables, including geopotential, humidity, temperature, and wind speed (Nguyen et al., 2023). The dataset extracted by LOTSA is based on a $64\times 128$ grid structure (Woo et al., 2024). In our experiments, we used data for year 2000 and kept a random subset of 500 samples out of the 8192 available grid points.
Oikolab Weather contains hourly measurements of eight meteorological variables from a weather station located near Monash University, Australia (Godahewa et al., 2021). All eight channels are kept in our experiments, treating them as univariate samples (channel independence).
Jena Weather contains 21 meteorological indicators, such as air temperature, humidity, etc. collected in 2020 at a 10-minute sampling rate from a weather station in Germany. Sourced by Wu et al. (2021) from https://www.bgc-jena.mpg.de/wetter/. All 21 variables are kept in our experiments, treating them as univariate samples (channel independence).
Weather contains hourly measurements of 11 meteorological variables (including temperatures, wind speed and direction, humidity, altimeter) in the US, during the period 2010-2013. We used the 11 variables in our experiments, handling them with channel independence. Sourced by Zhou et al. (2021) from https://www.ncei.noaa.gov/data/local-climatological-data/.
Transport domain
Pedestrian Counts* contains hourly pedestrian counts captured from 66 sensors in Melbourne city starting from May 2009 and up to 2020. It is part of the Monash Time Series Forecasting Library (Godahewa et al., 2021) and is sourced from the City of Melbourne.
Traffic is a collection of 48 months (2015-2016) road occupancy data from the California Department of Transportation (Lai et al., 2018). The road occupancy rates (between 0 and 1) are measured by different sensors on San Francisco Bay area freeways.
Loop Seattle contains one-year traffic state data from 323 sensor stations in the Greater Seattle Area, in 2015. Data were collected from inductive loop detectors deployed on four connected freeways (Cui et al., 2019).
PEMS03 is a highway traffic dataset collected by Song et al. (2020) from the California Department of Transportation Performance Measurement System (PeMS). PEMS03 contains 358 sensors with measurements from January to November 2018. We used four weeks of data for validation, four weeks for testing and the first 35 days for training.
PEMS BAY contains six months of measurements of traffic speed from 325 sensors in the Bay Area, California, in 2017 (Li et al., 2018). The train / validation / test split is 0.65 / 0.155 (28 days) / 0.195.
SHMETRO contains the passengers inflow and outflow measurements of 288 subway stations in Shanghai for three months in 2016 (Liu et al., 2020). We kept four weeks of data for validation, four weeks for testing and the first 35 days for training.
M Dense contains measurements of traffic intensity (number of cars per hour) from 30 sensors located in the city of Madrid, Spain, in 2018-2019. The dataset was sourced by the LibCity library through the open data portal of the Municipality of Madrid (de Medrano & Aznarte, 2020; Jiang et al., 2023).
B.2 Datasets with covariates
Table˜4 shows the key properties of the three datasets used for the evaluation of imputation with additional covariates. Each dataset is split chronologically intro train, validation and test splits with respective fractions 0.7, 0.1 and 0.2. Four-week segments are generated in the same manner as in Section˜B.1 for the univariate experiments.
Table 4: All datasets with covariates used in our experiments and their key properties.
Dataset Release Freq Target Covariate Series Num. Test
Platform Series Length Segments
PV-France RTE, 1H 1 1 8760 38
Meteo France
Wind-France RTE, 1H 1 1 8760 38
Meteo France
Load-France RTE, 30min 1 1 17520 41
Enedis
PV-France contains the aggregated photovoltaic (PV) power production (target variable) and the average solar irradiance (covariate) in the southern French region Occitanie in 2021. This dataset was obtained by aggregating two sources of data. (i) The target PV power production is provided by France’s Transmission System Operator (RTE) and extracted through their data portal https://www.rte-france.com/en/eco2mix/download-indicators. (ii) The global solar irradiance is obtained from the French weather institute Meteo France https://meteo.data.gouv.fr/datasets/donnees-climatologiques-de-base-horaires/. We aggregated the in-situ observations at the department level into a region-level irradiance.
Wind-France contains the aggregated wind power production (target variable) and the wind speed (covariate) in the northern French region Hauts-de-France in 2021. Similarly to PV-France, this dataset is obtained respectively via RTE’s data portal for the wind power production and via Meteo France for the wind speed.
Load-France contains the total French electricity demand (target variable) and the average temperature (covariate) in 2022. Similarly to PV-France and Wind-France, the total electricity demand is obtained from RTE’s data portal. The national temperature is provided by Enedis, the French distribution grid operator, from https://data.enedis.fr/explore/dataset/donnees-de-temperature-et-de-pseudo-rayonnement/information/.
Appendix C Univariate benchmark: extensive results
C.1 Full results
Section˜C.1 reports the complete univariate imputation benchmark across all datasets and missingness settings. These detailed results complement the main text by providing per-dataset performance for every method, allowing a finer comparison of their robustness and consistency across different missingness patterns and time series domain. For more details about the datasets please refer to Section˜B.1. All experiments were carried out on single NVIDIA A100-40G or H100-80G GPUs.
Complete univariate benchmark results for all datasets and missingness settings. Best in bold, second-best underlined. Settings Pointwise 1 and Pointwise 2: respectively 50% and 70% of observations removed at random. Settings Blocks 1 and Blocks 2: respectively two and four entire days removed at random.
Foundation models Task Specific Models Local Models
Dataset Setting
TabPFN
-TS
MoTM
SAITS
BRITS
CSDI
Time
mixerpp
Times
net
TSla
net
Linear
Seasonal
Naive
LOCF
[…]
Foundation models Task Specific Models Local Models
Dataset Setting
TabPFN
-TS
MoTM
SAITS
BRITS
CSDI
Time
mixerpp
Times
net
TSla
net
Linear
Seasonal
Naive
LOCF
[…]
Pointwise 2 0.223 0.240 0.309 0.432 0.284 0.825 0.830 0.833 0.330 0.587 0.545
Blocks 1 0.272 0.332 0.399 0.445 0.387 0.829 0.817 0.831 0.857 0.478 0.904
Blocks 2 0.280 0.336 0.405 0.471 0.395 0.829 0.819 0.832 0.861 0.473 0.904
BDG2-Rat Pointwise 1 0.196 0.231 0.266 0.279 0.251 0.813 0.812 0.818 0.247 0.587 0.380
Pointwise 2 0.256 0.273 0.339 0.338 0.314 0.813 0.814 0.818 0.334 0.646 0.507
Blocks 1 0.349 0.400 0.495 0.503 0.494 0.813 0.808 0.816 0.743 0.536 0.811
Blocks 2 0.355 0.402 0.497 0.508 0.498 0.813 0.808 0.816 0.744 0.536 0.814
Borealis Pointwise 1 0.417 0.519 0.403 0.407 0.687 0.598 0.535 0.596 0.442 0.647 0.543
Pointwise 2 0.488 0.583 0.468 0.442 0.672 0.594 0.545 0.597 0.505 0.674 0.601
Blocks 1 0.536 0.646 0.518 0.508 0.685 0.627 0.540 0.612 0.633 0.662 0.718
Blocks 2 0.522 0.629 0.519 0.504 0.658 0.612 0.537 0.602 0.658 0.612 0.638
Covid19 Energy Pointwise 1 0.075 0.099 0.399 0.307 1.113 0.405 0.414 0.852 0.163 0.438 0.387
Pointwise 2 0.132 0.127 0.417 0.436 1.118 0.486 0.417 0.854 0.292 0.483 0.570
Blocks 1 0.202 0.222 0.432 0.629 1.117 0.604 0.440 0.858 0.949 0.399 0.975
Blocks 2 0.201 0.232 0.436 0.638 1.114 0.599 0.461 0.842 0.928 0.392 0.939
GFC12 Load Pointwise 1 0.143 0.189 0.188 0.263 0.174 0.789 0.776 0.803 0.232 0.603 0.448
Pointwise 2 0.237 0.231 0.280 0.457 0.248 0.789 0.787 0.804 0.370 0.670 0.600
Blocks 1 0.353 0.382 0.417 0.524 0.428 0.793 0.788 0.797 0.810 0.546 0.868
Blocks 2 0.363 0.387 0.425 0.541 0.436 0.799 0.796 0.803 0.814 0.546 0.871
Hog Pointwise 1 0.196 0.240 0.245 0.244 0.279 0.585 0.785 0.802 0.216 0.701 0.322
Pointwise 2 0.260 0.286 0.320 0.326 0.350 0.593 0.795 0.801 0.280 0.759 0.418
Blocks 1 0.396 0.458 0.518 0.562 0.569 0.651 0.766 0.799 0.555 0.640 0.668
Blocks 2 0.406 0.464 0.528 0.568 0.580 0.655 0.766 0.795 0.564 0.649 0.673
Ideal Pointwise 1 0.501 0.570 0.443 0.464 0.603 0.531 0.702 0.691 0.526 0.678 0.620
Pointwise 2 0.571 0.644 0.473 0.504 0.638 0.572 0.706 0.692 0.592 0.701 0.683
Blocks 1 0.558 0.667 0.498 0.543 0.655 0.531 0.706 0.696 0.730 0.650 0.724
Blocks 2 0.564 0.658 0.485 0.522 0.648 0.497 0.688 0.676 0.699 0.657 0.744
PDB Pointwise 1 0.062 0.094 0.337 0.408 1.122 0.449 0.395 0.864 0.191 0.375 0.435
Pointwise 2 0.119 0.121 0.373 0.556 1.128 0.550 0.599 0.858 0.338 0.413 0.631
Blocks 1 0.171 0.197 0.353 0.637 1.108 0.726 0.265 0.832 0.982 0.331 1.004
Blocks 2 0.177 0.201 0.356 0.653 1.113 0.740 0.291 0.847 1.006 0.336 1.010
KDD Cup2022 Pointwise 1 0.099 0.237 0.115 0.764 0.171 0.764 0.758 0.775 0.092 0.955 0.139
Pointwise 2 0.120 0.255 0.161 0.768 0.217 0.767 0.766 0.776 0.115 0.971 0.176
Blocks 1 0.571 0.636 0.628 0.763 0.914 0.763 0.764 0.769 0.467 0.939 0.654
Blocks 2 0.566 0.634 0.624 0.758 0.912 0.758 0.761 0.765 0.464 0.927 0.660
ERA5 geo. Pointwise 1 0.091 0.168 0.149 0.161 0.163 0.251 0.813 0.815 0.104 0.728 0.224
Pointwise 2 0.146 0.208 0.219 0.275 0.253 0.427 0.814 0.815 0.162 0.791 0.322
Blocks 1 0.333 0.452 0.525 0.905 0.680 0.709 0.806 0.808 0.473 0.681 0.627
Blocks 2 0.336 0.449 0.524 0.908 0.680 0.710 0.807 0.809 0.475 0.681 0.629
ERA5 humidity Pointwise 1 0.129 0.212 0.169 0.177 0.216 0.413 0.788 1.670 0.131 0.664 0.240
Pointwise 2 0.200 0.254 0.233 0.270 0.315 0.544 0.792 1.320 0.191 0.748 0.314
Blocks 1 0.336 0.434 0.415 0.501 0.608 0.558 0.787 2.059 0.372 0.603 0.500
Blocks 2 0.337 0.431 0.417 0.503 0.608 0.559 0.784 1.989 0.373 0.606 0.506
ERA5 temp. Pointwise 1 0.094 0.168 0.158 0.157 0.177 0.481 0.817 1.084 0.112 0.700 0.237
Pointwise 2 0.150 0.208 0.231 0.250 0.261 0.603 0.815 0.947 0.174 0.765 0.340
Blocks 1 0.327 0.438 0.520 0.534 0.677 0.725 0.810 1.292 0.504 0.651 0.647
Blocks 2 0.331 0.436 0.521 0.539 0.681 0.725 0.808 1.258 0.503 0.653 0.647
ERA5 wind Pointwise 1 0.124 0.225 0.176 0.182 0.195 0.797 0.790 3.795 0.127 0.945 0.252
Pointwise 2 0.201 0.281 0.248 0.298 0.285 0.793 0.795 2.637 0.187 0.994 0.346
Blocks 1 0.461 0.604 0.552 0.706 0.722 0.802 0.800 5.137 0.465 0.902 0.685
Blocks 2 0.464 0.604 0.553 0.713 0.720 0.799 0.797 4.954 0.467 0.910 0.687
Oikolab Weather Pointwise 1 0.150 0.227 0.207 0.207 0.226 0.802 0.824 0.827 0.164 0.811 0.313
Pointwise 2 0.228 0.278 0.294 0.327 0.321 0.805 0.826 0.828 0.248 0.857 0.436
Blocks 1 0.449 0.529 0.581 0.669 0.739 0.820 0.815 0.823 0.638 0.766 0.786
Blocks 2 0.454 0.534 0.584 0.674 0.735 0.821 0.818 0.826 0.636 0.770 0.787
Pedestrian Counts Pointwise 1 0.150 0.196 0.240 0.268 0.176 0.786 0.804 0.812 0.329 0.347 0.526
Pointwise 2 0.200 0.239 0.289 0.387 0.213 0.793 0.806 0.812 0.447 0.383 0.684
Blocks 1 0.172 0.254 0.243 0.619 0.210 0.797 0.795 0.810 1.001 0.307 1.002
Blocks 2 0.178 0.260 0.252 0.632 0.214 0.797 0.797 0.811 1.000 0.310 0.997
Traffic Pointwise 1 0.172 0.237 0.174 0.208 0.208 0.735 0.735 0.744 0.288 0.380 0.498
Pointwise 2 0.216 0.285 0.228 0.242 0.242 0.738 0.738 0.744 0.421 0.416 0.670
Blocks 1 0.204 0.301 0.235 0.252 0.252 0.731 0.731 0.744 0.985 0.340 0.985
Blocks 2 0.210 0.307 0.242 0.256 0.256 0.732 0.732 0.744 0.986 0.341 0.985
SHMETRO Pointwise 1 0.123 0.297 0.530 0.299 1.413 0.642 0.314 0.675 0.168 0.254 0.275
Pointwise 2 0.143 0.325 0.537 0.376 1.493 0.641 0.450 0.673 0.219 0.280 0.363
Blocks 1 0.190 0.407 0.538 0.632 1.283 0.642 0.277 0.671 0.963 0.217 0.970
Blocks 2 0.197 0.409 0.546 0.632 1.441 0.642 0.289 0.670 0.917 0.220 0.910
PEMS03 Pointwise 1 0.133 0.189 0.353 0.360 1.300 0.862 0.439 0.883 0.148 0.349 0.183
Pointwise 2 0.148 0.197 0.351 0.496 1.440 0.862 0.605 0.881 0.156 0.377 0.206
Blocks 1 0.233 0.333 0.376 0.862 1.150 0.867 0.323 0.879 1.098 0.330 1.109
Blocks 2 0.229 0.332 0.367 0.855 1.176 0.861 0.312 0.872 1.083 0.315 1.097
PEMS BAY Pointwise 1 0.127 0.306 0.188 0.195 0.863 0.558 0.595 0.642 0.121 0.458 0.170
Pointwise 2 0.146 0.322 0.197 0.203 0.851 0.558 0.615 0.642 0.145 0.490 0.207
Blocks 1 0.324 0.513 0.423 0.435 0.865 0.555 0.638 0.641 0.775 0.422 0.800
Blocks 2 0.329 0.521 0.423 0.436 0.875 0.560 0.639 0.643 0.776 0.427 0.795
ETT1-15T Pointwise 1 0.180 0.288 0.227 0.260 0.267 0.535 0.500 0.793 0.183 0.610 0.256
Pointwise 2 0.208 0.311 0.262 0.357 0.322 0.585 0.619 0.793 0.213 0.648 0.316
Blocks 1 0.445 0.504 0.662 0.760 0.688 0.783 0.626 0.791 0.821 0.584 0.890
Blocks 2 0.452 0.513 0.660 0.759 0.698 0.780 0.648 0.790 0.814 0.581 0.883
ETT1-1H Pointwise 1 0.230 0.279 0.314 0.309 0.380 0.787 0.596 0.797 0.279 0.588 0.446
Pointwise 2 0.306 0.326 0.404 0.399 0.482 0.784 0.692 0.797 0.381 0.630 0.582
Blocks 1 0.412 0.465 0.577 0.603 0.683 0.793 0.610 0.796 0.820 0.562 0.882
Blocks 2 0.419 0.473 0.583 0.617 0.690 0.793 0.617 0.796 0.828 0.557 0.891
ETT2-1H Pointwise 1 0.281 0.321 0.333 0.341 0.450 0.778 0.758 0.801 0.293 0.698 0.426
Pointwise 2 0.350 0.371 0.395 0.420 0.514 0.780 0.761 0.801 0.365 0.750 0.525
Blocks 1 0.480 0.526 0.568 0.635 0.697 0.790 0.781 0.798 0.686 0.674 0.784
Blocks 2 0.490 0.527 0.575 0.642 0.692 0.788 0.783 0.797 0.689 0.663 0.785
ETT2-15T Pointwise 1 0.215 0.332 0.312 0.275 0.308 0.768 0.717 0.801 0.216 0.716 0.280
Pointwise 2 0.247 0.356 0.339 0.351 0.368 0.768 0.781 0.801 0.249 0.764 0.334
Blocks 1 0.543 0.554 0.592 0.756 0.713 0.792 0.777 0.799 0.675 0.688 0.765
Blocks 2 0.547 0.553 0.594 0.759 0.718 0.797 0.783 0.805 0.697 0.691 0.785
Solar-1H Pointwise 1 0.132 0.131 0.141 0.297 0.140 0.304 0.632 0.828 0.216 0.260 0.407
Pointwise 2 0.176 0.168 0.213 0.415 0.187 0.310 0.685 0.828 0.365 0.277 0.589
Blocks 1 0.219 0.228 0.237 0.347 0.264 0.299 0.534 0.829 0.930 0.239 0.921
Blocks 2 0.219 0.229 0.236 0.357 0.265 0.298 0.550 0.830 0.932 0.239 0.924
Jena Weather 10T Pointwise 1 0.086 0.190 0.177 0.132 0.174 0.724 0.339 0.748 0.080 0.635 0.123
Pointwise 2 0.101 0.207 0.196 0.188 0.201 0.725 0.490 0.748 0.096 0.688 0.157
Blocks 1 0.378 0.453 0.557 0.681 0.685 0.739 0.699 0.746 0.527 0.591 0.640
Blocks 2 0.384 0.459 0.561 0.681 0.712 0.736 0.700 0.743 0.544 0.598 0.644
Jena Weather 1H Pointwise 1 0.156 0.224 0.346 0.223 0.323 0.738 0.675 0.754 0.170 0.638 0.293
Pointwise 2 0.214 0.272 0.394 0.316 0.418 0.741 0.725 0.754 0.243 0.700 0.394
Blocks 1 0.366 0.463 0.532 0.583 0.674 0.736 0.703 0.744 0.552 0.608 0.662
Blocks 2 0.370 0.464 0.532 0.591 0.679 0.744 0.711 0.750 0.555 0.606 0.653
Loop Seattle 5T Pointwise 1 0.233 0.371 0.326 0.345 0.309 0.661 0.708 0.721 0.246 0.644 0.300
Pointwise 2 0.261 0.387 0.345 0.368 0.334 0.661 0.708 0.721 0.266 0.672 0.331
Blocks 1 0.469 0.590 0.596 0.665 0.573 0.662 0.709 0.722 0.842 0.611 0.895
Blocks 2 0.474 0.597 0.600 0.675 0.574 0.662 0.710 0.722 0.842 0.612 0.898
Loop Seattle 1H Pointwise 1 0.279 0.337 0.321 0.402 0.325 0.505 0.715 0.721 0.379 0.576 0.530
Pointwise 2 0.350 0.394 0.376 0.502 0.373 0.524 0.713 0.721 0.486 0.615 0.653
Blocks 1 0.345 0.417 0.367 0.511 0.394 0.505 0.715 0.720 0.845 0.532 0.884
Blocks 2 0.354 0.425 0.376 0.523 0.400 0.511 0.717 0.721 0.851 0.533 0.890
M Dense Pointwise 1 0.209 0.233 0.274 0.422 0.276 0.851 0.456 0.855 0.341 0.447 0.552
Pointwise 2 0.252 0.276 0.315 0.604 0.316 0.854 0.634 0.855 0.473 0.484 0.722
Blocks 1 0.225 0.284 0.288 0.415 0.302 0.849 0.433 0.856 1.039 0.405 1.060
Blocks 2 0.228 0.288 0.291 0.433 0.305 0.847 0.455 0.854 1.041 0.407 1.061
Enedis LDM Small Pointwise 1 0.340 0.480 0.299 0.317 0.356 0.499 0.584 0.603 0.340 0.373 0.439
Pointwise 2 0.440 0.534 0.338 0.358 0.415 0.499 0.589 0.594 0.419 0.396 0.524
Blocks 1 0.358 0.748 0.337 0.358 0.451 0.496 0.581 0.637 0.647 0.351 0.719
Blocks 2 0.366 0.515 0.341 0.362 0.456 0.498 0.581 0.633 0.707 0.345 0.722
London Small Pointwise 1 0.439 0.573 0.432 0.434 0.837 0.596 0.650 0.965 0.480 0.666 0.573
Pointwise 2 0.490 0.632 0.465 0.483 0.843 0.596 0.651 0.760 0.534 0.679 0.638
Blocks 1 0.521 0.622 0.537 0.571 0.822 0.595 0.650 1.631 0.794 0.656 0.836
Blocks 2 0.528 0.628 0.512 0.576 0.825 0.595 0.650 1.570 0.798 0.656 0.836
Spanish Energy Pointwise 1 0.131 0.200 0.200 0.235 1.716 0.782 0.812 0.802 0.164 0.680 0.298
Pointwise 2 0.207 0.246 0.290 0.346 1.905 0.783 0.810 0.802 0.253 0.735 0.413
Blocks 1 0.400 0.472 0.559 0.633 1.367 0.787 0.788 0.801 0.614 0.642 0.719
Blocks 2 0.399 0.469 0.559 0.637 1.384 0.784 0.788 0.800 0.609 0.635 0.715
Weather Pointwise 1 0.257 0.298 0.290 0.360 3.369 0.801 0.804 0.804 0.259 0.739 0.373
Pointwise 2 0.311 0.342 0.364 0.458 3.928 0.792 0.803 0.804 0.322 0.803 0.472
Blocks 1 0.456 0.494 0.592 0.689 1.680 0.803 0.797 0.800 0.620 0.686 0.736
Blocks 2 0.455 0.488 0.590 0.691 1.909 0.806 0.799 0.801 0.618 0.682 0.735
C.2 Performance breakdown by missingness pattern
(a) 50 % pointwise missing values
(b) 70 % pointwise missing values
(c) Two days missing (missing blocks)
(d) Four days missing (missing blocks)
Figure 7: Normalized MAEs results for each distinct setting of missingness patterns.
Figure˜7 provides a detailed breakdown of the univariate benchmark results presented in the main paper, focusing on the four distinct missingness settings introduced in Section˜3.1. Specifically, we report aggregated z-normalized MAE scores for (a) 50% and (b) 70% pointwise missingness, as well as for (c) two-day and (d) four-day block missingness. These complementary analyses aim to highlight the robustness and consistency of model performance across different temporal corruption patterns.
Results.
The per-setting breakdown shows a consistent advantage of TabPFN-TS across all four settings. However, the relative ranking of other methods varies with the missingness pattern: in the first pointwise setting (50%) Linear is actually stronger than MoTM and remains roughly tied with MoTM at 70% pointwise, indicating that simple local interpolation can excel when gaps are sparse. By contrast, in block-missing scenarios (two and four days) MoTM substantially outperforms Linear, showing superior robustness to structured, long gaps. SAITS is the best supervised baseline but is generally behind the foundation models; its relative position is closer to MoTM in some settings (e.g. pointwise). Overall, these results emphasize (i) the stability of TabPFN-TS across patterns, (ii) that local methods can be competitive for sparse pointwise missingness, and (iii) that foundation/zero-shot models better handle large or by-block gaps.
C.3 Quantile predictions
In addition to pointwise imputations, we evaluate the probabilistic capabilities of the time-indexed foundation models by assessing the quality of their predicted quantiles. This analysis complements the deterministic metrics by measuring how well models can represent uncertainty around missing points. Accurate quantile estimation is particularly important for time-series imputation under distributional shifts or irregular sampling, as it reflects the model’s ability to produce calibrated and reliable uncertainty quantification.
C.3.1 Implementation details
Weighted Quantile Loss (WQL) definition.
We evaluate probabilistic imputations using the Weighted Quantile Loss (WQL), introduced by Koenker & Hallock (2001) and adopted in probabilistic forecasting benchmarks such as Gneiting et al. (2007); Gasthaus et al. (2019). Given a predicted $\alpha$-quantile $q$ of an observation $x$, the quantile loss is defined as:
$\mathrm{QL}_{\alpha}(q,x)=\begin{cases}\alpha(x-q),&\text{if }x>q,\\ (1-\alpha)(q-x),&\text{otherwise.}\end{cases}$ (1)
To aggregate this quantity across all series and time steps, we compute a weighted average normalized by the absolute scale of targets:
$\mathrm{WQL}_{\alpha}=\frac{2\sum_{i,t}\mathrm{QL}_{\alpha}(q^{(\alpha)}_{i,t},x_{i,t})}{\sum_{i,t}|x_{i,t}|}.$ (2)
We then average over a finite set of quantile levels $\{\alpha_{1},\ldots,\alpha_{K}\}$:
$\mathrm{WQL}=\frac{1}{K}\sum_{j=1}^{K}\mathrm{WQL}_{\alpha_{j}}.$ (3)
Quantiles are evaluated at $\alpha\in\{0.1,0.2,\ldots,0.9\}$ ($K=9$). Being a weighted average of quantile losses across levels, WQL approximates the Continuous Ranked Probability Score (CRPS), a standard metric for probabilistic accuracy.
CSDI and SAITS for quantile evaluation.
For CSDI (a stochastic diffusion imputer), we estimate quantiles post hoc by Monte Carlo sampling at test time. Conditioned on the observed context and the evaluation mask, we draw $S$ imputations $\{\tilde{x}^{(s)}_{i,t}\}_{s=1}^{S}$ from the model’s posterior and compute empirical quantiles per index $(i,t)$: $\hat{q}^{(\alpha)}_{i,t}\;=\;\operatorname{Quantile}_{\alpha}\!\left(\{\tilde{x}^{(s)}_{i,t}\}_{s=1}^{S}\right).$ We use $S=50$ across all datasets. For SAITS, we train one independent model per quantile level $\alpha\in\{0.1,\dots,0.9\}$. Each model uses the standard pinball (quantile) loss computed only on the evaluation mask, with the same masking patterns as in the pointwise setting. At inference, we obtain the full set of quantiles by stacking the predictions from the $K$ independently trained models. Both models are evaluated on identical imputation splits and masking ratios; WQL is computed only on masked targets and normalized by the absolute scale as described in Eq. (2).
C.3.2 Quantitative results
Setting.
We evaluate the models for uncertainty quantification on 11 univariate datasets, namely: BDG2-Bear & Rat, Covid19 Energy, GFC12 Load, Hog, Jena Weather 10T, Jena Weather 1H, Oikolab Weather, PDB, Pedestrian Counts and Weather, described in Appendix˜B. All models are trained with the same masking ratios and evaluated on the same imputation splits to ensure comparability.
Table 5: Complete WQL results on 11 datasets. For each setting, the best score (lowest) is in bold and the second best is underlined.
Dataset Setting
TabPFN-TS
MoTM
CSDI
SAITS-adapted
BDG2-Bear Pointwise 1 0.165 0.208 0.183 0.294
Pointwise 2 0.208 0.256 0.223 0.373
Blocks 1 0.253 0.354 0.336 0.493
Blocks 2 0.257 0.355 0.338 0.493
BDG2-Rat Pointwise 1 0.191 0.239 0.194 0.321
Pointwise 2 0.203 0.256 0.242 0.412
Blocks 1 0.338 0.436 0.387 0.617
Blocks 2 0.334 0.431 0.382 0.612
Covid19 Energy Pointwise 1 0.071 0.097 0.758 0.353
Pointwise 2 0.124 0.130 0.757 0.418
Blocks 1 0.187 0.236 0.767 0.522
Blocks 2 0.187 0.246 0.759 0.521
GFC12 Load Pointwise 1 0.142 0.192 0.148 0.258
Pointwise 2 0.212 0.249 0.233 0.371
Blocks 1 0.349 0.427 0.383 0.528
Blocks 2 0.349 0.423 0.381 0.523
Hog Pointwise 1 0.196 0.249 0.230 0.322
Pointwise 2 0.254 0.311 0.289 0.410
Blocks 1 0.395 0.513 0.505 0.698
Blocks 2 0.392 0.507 0.504 0.695
Jena Weather 10T Pointwise 1 0.113 0.225 0.141 0.418
Pointwise 2 0.134 0.235 0.164 0.445
Blocks 1 0.397 0.508 0.706 0.791
Blocks 2 0.394 0.500 0.712 0.781
Jena Weather 1H Pointwise 1 0.181 0.254 0.246 0.354
Pointwise 2 0.234 0.313 0.317 0.437
Blocks 1 0.384 0.548 0.619 0.697
Blocks 2 0.382 0.528 0.605 0.693
Oikolab Weather Pointwise 1 0.145 0.225 0.157 0.248
Pointwise 2 0.212 0.291 0.234 0.358
Blocks 1 0.431 0.584 0.578 0.761
Blocks 2 0.420 0.572 0.568 0.748
PDB Pointwise 1 0.057 0.084 0.766 0.272
Pointwise 2 0.111 0.118 0.764 0.347
Blocks 1 0.170 0.209 0.774 0.391
Blocks 2 0.171 0.205 0.773 0.406
Pedestrian Counts Pointwise 1 0.147 0.201 0.171 0.228
Pointwise 2 0.189 0.253 0.221 0.300
Blocks 1 0.170 0.260 0.199 0.277
Blocks 2 0.173 0.264 0.203 0.279
Weather Pointwise 1 0.252 0.301 0.263 0.378
Pointwise 2 0.303 0.356 0.322 0.456
Blocks 1 0.448 0.528 0.601 0.734
Blocks 2 0.431 0.510 0.582 0.716
Results.
Table˜5 reports the WQL scores across all eleven datasets and four missingness settings. TabPFN-TS achieves the lowest WQL across most datasets, showing its ability to produce both accurate and well-calibrated uncertainty estimates. MoTM follows with higher WQL but consistent behavior across datasets and settings. CSDI ranks third overall, with performances illustrating its limited generalization abilities: (i) consistently with its training procedure, good scores are obtained in the pointwise settings; (ii) except on small-size training sets, such as Covid19 Energy or PDB; (iii) CSDIgenerally suffers from significantly higher losses on the unseen Block settings. Finally, the SAITS variant adapted for quantile prediction yields the highest WQL in most settings, highlighting the shortcomings of this straightforward per-quantile adaptation compared with models expressly designed for probabilistic imputation.
C.3.3 Qualitative results
In Figure˜8 and Figure˜9, we present qualitative examples of quantile-based imputations for both TabPFN-TS and MoTM on segments where 70% of the values were removed in a pointwise manner.
For each plot, we display the median prediction together with the inter-quantile ranges $[5,95]$ and $[25,75]$, which highlight different levels of predictive uncertainty.
Results.
The visualizations in Figure˜8 and Figure˜9 confirm that both TabPFN-TS and MoTM provide effective quantile-based imputations, successfully reconstructing the signal while providing meaningful uncertainty estimates. TabPFN-TS particularly excels at generating high-fidelity reconstructions that closely follow the ground truth, capturing fine-grained temporal details and high-frequency oscillations. Its strength lies in its highly adaptive uncertainty quantification; on the volatile Borealis dataset, the quantile ranges adeptly widen to reflect increased predictive uncertainty around sharp peaks, demonstrating a sophisticated understanding of local signal dynamics. MoTM, for its part, also delivers strong performance by producing robust, albeit smoother, imputations that effectively capture the main trends and cyclical patterns in datasets like Hog and Era5. Its more regular uncertainty bands provide a consistent and reliable confidence envelope around the reconstructed signal. In summary, while both models prove highly competent for this task, they exhibit different strengths: TabPFN-TS favors a detailed, high-fidelity reconstruction, whereas MoTM prioritizes capturing the underlying trend with stable uncertainty.
(a) Jena Weather dataset
(b) Borealis dataset
Figure 8: TabPFN-TS qualitative quantile results in the 70% missing values scenario (Pointwise 2).
(a) Hog dataset
(b) ERA5 geopotential dataset
Figure 9: MoTM qualitative quantile results in the 70% missing values scenario (Pointwise 2).
C.4 Experiments on lower sampling rates datasets
The experiments presented in Section˜3.1 focus on datasets with relatively high temporal resolutions (5min, 10min, 15min, 30min, 1h). In this section, we investigate whether time-index foundation models can generalize to significantly lower sampling rates, such as daily or weekly observations. This setup evaluates their robustness to long-term dependencies and coarser temporal granularity, which are common in macroeconomic, energy, or demographic data.
Datasets.
We consider four publicly available low-frequency datasets: Births Daily, M4 Daily, Births Weekly, and M4 Weekly, with statistics summarized in Table˜6. They are part of the Monash Time Series Forecasting archive (Godahewa et al., 2021), and were downloaded from the GIFT-eval repository (Aksu et al., 2024). Each dataset exhibits distinct temporal behaviors—seasonality and periodicity are typically weaker at weekly scales, while daily data contain more regular cycles and higher variance. For each dataset, we apply the same four missingness regimes as in the main benchmark (two pointwise and two block-based scenarios), allowing a consistent comparison across frequencies.
Table 6: All datasets at lower sampling rates used in our experiments and their key properties.
Dataset Release Domain Freq Num. Series Num. Test
Platform Series Length Segments
Births Daily GIFT-eval Demo. 1D 1 3652 104
Births Weekly GIFT-eval Demo. 1W 1 1043 7
M4 Daily GIFT-eval Econ./Demo. 1D 2112 2954 80256
M4 Weekly GIFT-eval Econ./Demo 1W 172 947 860
Births Daily contains the daily number of births in the US between 1969 and 1988, as extracted from the R package mosaicData (Pruim et al., 2020). Births Weekly aggregates these statistics at a weekly frequency.
The M4 Forecasting Competition dataset contains a total of 100k time series, with six different frequencies and from diverse domains such as demography, macroeconomic, etc. (Makridakis et al., 2018; 2020). We used the subsets of, respectively, daily (M4 Daily) and weekly (M4 Weekly) time series.
Baselines and settings.
We evaluate a representative subset of imputation methods present in the main benchmark: local heuristics (Linear, Seasonal Naive), supervised models SAITS, CSDI, BRITS, and the two foundation models TabPFN-TS and MoTM. As before, TabPFN-TS and MoTM operate in a fully zero-shot setting, while SAITS is retrained for each dataset. All metrics are aggregated using the z-normalized Mean Absolute Error (MAE), and results are reported per dataset and missingness pattern in Table˜7.
Table 7: Complete MAE results on datasets with low sampling rates (daily or weekly). Best results are in bold, second-best are underlined.
Dataset Setting
TabPFN-TS
MoTM
SAITS
BRITS
CSDI
Linear
Seasonnal Naive
Births Daily Pointwise 1 0.281 0.359 0.804 0.324 1.049 0.880 0.286
Pointwise 2 0.340 0.409 0.814 0.454 1.095 0.941 0.318
Blocks 1 0.243 0.361 0.770 0.430 1.052 0.995 0.300
Blocks 2 0.272 0.361 0.783 0.447 1.064 0.988 0.291
M4 Daily Pointwise 1 0.182 0.248 0.185 0.211 0.223 0.159 0.507
Pointwise 2 0.246 0.304 0.234 0.288 0.273 0.195 0.606
Blocks 1 0.442 0.531 0.430 0.576 0.638 0.381 0.547
Blocks 2 0.454 0.534 0.441 0.591 0.614 0.385 0.559
Births Weekly Pointwise 1 0.293 0.337 0.896 0.748 11.999 0.332 0.492
Pointwise 2 0.335 0.358 0.917 0.778 14.635 0.385 0.602
Blocks 1 0.301 0.320 1.036 0.840 4.008 0.723 0.481
Blocks 2 0.324 0.338 0.802 0.713 5.458 0.762 0.502
M4 Weekly Pointwise 1 0.160 0.226 0.191 0.310 0.307 0.175 0.761
Pointwise 2 0.188 0.254 0.226 0.415 0.344 0.198 0.836
Blocks 1 0.244 0.326 0.430 0.635 0.435 0.318 0.643
Blocks 2 0.256 0.345 0.444 0.643 0.459 0.340 0.650
Average score 0.285 0.351 0.525 0.525 2.791 0.510 0.524
Results.
As shown in Table˜7, foundation models demonstrate robust performance on low-frequency data. TabPFN-TS achieves the best average score (0.285), followed by MoTM (0.351), confirming their ability to generalize to coarser temporal structures without retraining. Other baselines, including specialized models like SAITS and BRITS, are significantly outperformed and show no clear advantage over simpler methods, while CSDI struggles notably. Overall, these results highlight the superior generalization of our zero-shot models in low-frequency regimes compared to both classic baselines and other deep learning architectures.
Appendix D Experiments with covariates: extensive qualitative results
To better visualize the impact of incorporating covariates, we present qualitative results for both TabPFN-TS and MoTM. Figure˜10 and Figure˜11 show examples from three datasets (PV-France, Wind-France, and Load-France), illustrating how the inclusion of covariate information helps each model reconstruct the four one-day missing blocks more accurately.
(a) Wind-France
(b) PV-France
(c) Load-France
Figure 10: TabPFN-TS qualitative results with and without covariates in the four one-day missing blocks scenario (Blocks 2).
(a) Wind-France
(b) PV-France
(c) Load-France
Figure 11: MoTM qualitative results with and without covariates in the four one-day missing blocks scenario (Blocks 2).
Results.
As illustrated in Figure˜10 and 11, incorporating covariates visually improves the reconstructions for both TabPFN-TS and MoTM. The covariate-enhanced versions capture missing intervals more smoothly and better follow short-term temporal variations within the four one-day gaps. We observe that the interpolations of TabPFN-TS align more closely with the ground truth and preserve sharper transitions across missing regions. In contrast, if MoTM consistently benefits from covariate information, its reconstructions sometimes struggle to capture sudden changes with great accuracy. Overall, the qualitative plots confirm that the most visible gains are observed on the Wind-France and PV-France datasets, while the impact remains limited for Load-France. One might note that quantile predictions can also be generated in this covariate setting, in a similar fashion to the univariate examples presented in Figure˜11.